OpenStudy (anonymous):

Multiply&Simplify: (x^1/3 + x^-1/3)^2

OpenStudy (anonymous):

x^(2/3) + x^(-2/3) = x^(2/3) + 1/x^(2/3)

OpenStudy (anonymous):

Is it ^2/3 or ^2/9?

OpenStudy (anonymous):

just sub in values and check its the 2/3 one though

OpenStudy (anonymous):

Okay, How did you get the +1/x^(2/3)??

OpenStudy (anonymous):

when you have x^(-y), that is when some thing to the power of something negative, you can inverse that such that x^(-y) = 1/x^y

OpenStudy (anonymous):

j_sims22, if you multiply the original "x^-1/3" by the exponent "2", you get "x^-2/3". This is the same as 1/x^2/3.

OpenStudy (anonymous):

hipster.. where did u get ur 2??

OpenStudy (anonymous):

Thank you both for helping me :)